OpenStudy (anonymous):

how do i write the equation in slope-intercept form, and then find the slope and y-intercept for 2x-y=-15

OpenStudy (triciaal):

rewrite to let the subject be y y = ?

OpenStudy (anonymous):

I'm supposed to isolate Y correct?

OpenStudy (triciaal):

yes

OpenStudy (triciaal):

the slope will be the coefficient of x and the intercept is the constant

OpenStudy (anonymous):

ok so 2x-y=-15 -y=-2x-15 y=2x+15

OpenStudy (triciaal):

another way to see the intercept is to ask what is the value of y when x = 0

OpenStudy (triciaal):

yes you got it

OpenStudy (anonymous):

lol srry I'm having a hard time I'm middle school

OpenStudy (anonymous):

ok so how do i write it in slope intercept form..or did i do that

OpenStudy (triciaal):

don't be sorry its ok here to help it takes practice to master the concepts

OpenStudy (triciaal):

that is slope intercept form y = mx + b

OpenStudy (triciaal):

slope m intercept b

OpenStudy (anonymous):

OH so did find the y intercept

OpenStudy (anonymous):

Thank you

OpenStudy (triciaal):

yes y intercept means what is the value of y when x = 0

OpenStudy (anonymous):

so the slope intercept form is y=2x+15? how do i find the slope? how do i find the y-intercept?

OpenStudy (triciaal):

you are welcome. keep practicing and it will get easier. your high school will be a breeze if you understand the terms in middle school

OpenStudy (triciaal):

y = mx + b the slope is m the intercept is b now match them up

OpenStudy (triciaal):

y = mx + b y = 2x + 15 is it clear now?

OpenStudy (anonymous):

OOOHHHHHHHHH yea i get it thank you.

OpenStudy (anonymous):

ill see you later thank you